Mohr Company purchases a machine at the beginning of the year at a cost of $26,000. The machine is depreciated using the double-declining-balance method. The machine’s useful life is estimated to be 5 years with a $5,000 salvage value. Depreciation expense in year 2 is:

Multiple Choice

$15,600.

$10,400.

$8,400.

$6,240.

$5,200.
Answer #1

Answer

  • Correct Answer = Option #4: $ 6,240 = Depreciation expense for Year 2
  • Working

A

Cost

$            26,000.00

B

Residual Value

$              5,000.00

C=A - B

Depreciable base

$            21,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$              4,200.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$              26,000.00

40.00%

$         10,400.00

$             15,600.00

2

$              15,600.00

40.00%

$            6,240.00 = Answer

$                9,360.00